OpenStudy (anonymous):

Simplify completely 40v^2 20V^3 ----- ÷ ----- = ? 35v^4 5v

OpenStudy (anonymous):

cancel the terms with 5 and the poewrs accordingly

OpenStudy (anonymous):

so cancel everything by 5 and v

OpenStudy (anonymous):

@shkrina

OpenStudy (jhannybean):

No,can't cancel until you make a common denominator.

OpenStudy (anonymous):

examples please

OpenStudy (anonymous):

so is it 2 ---- 7v^4

OpenStudy (anonymous):

can you teach me ? lol @shkrina whats thre common denominator ? 7v^2 ?

OpenStudy (jhannybean):

\[\large \frac{40v^2}{35v^4}+\frac{20v^3}{5v}\] make a common denominator by multiplying the second fraction with 7v^3 \[\large \frac{40v^2}{35v^4}+\frac{140v^6}{35v^4}\]combine into one fraction \[\large \frac{40v^2+140v^6}{35v^4}\]factor out a 20v^2 from numerator \[\large \frac{20v^2(20 + 7v^4)}{35v^4}\]Now simplify \[\large \frac{4(20 + 7v^4)}{7v^2}\]

OpenStudy (anonymous):

@Jhannybean have explained hope understand that ..... ..hope u r cleared....

OpenStudy (jhannybean):

Now distribute the 4. \[\large \frac{80+28v^4}{7v^2}\] divide each term in the numerator with the denominator. \[\large \frac{80}{7v^2}+4v^2\]

OpenStudy (anonymous):

so its 80+28v^4 --------- 7v^2

OpenStudy (jhannybean):

yep.

OpenStudy (anonymous):

now i ?

OpenStudy (anonymous):

@Jhannybean i hope its 8 not 80

OpenStudy (anonymous):

i think it is 8...

OpenStudy (jhannybean):

Ohyes you are right, Sorry about that.\[\large \frac{20v^2(2 + 7v^4)}{35v^4}\]

OpenStudy (anonymous):

then itll be 2 ---- 7v^4 ?

OpenStudy (jhannybean):

\[\large \frac{4(2+7v^4)}{7v^2}\]

OpenStudy (anonymous):

insted of doing this long process u can directly cancel by 5 and the powers u will get the same ans now do LCM

OpenStudy (anonymous):

8+28v^4 --------- 7v^2

OpenStudy (jhannybean):

\[\large \frac{8+28v^4}{7v^2}\] But it's not the proper way, you can't cancel stuff out in the fractions unless you factor stuff out.

OpenStudy (anonymous):

so now what ? these are the answer choices btw 2/7v^4 7/32 7v^4/32 32/7

OpenStudy (anonymous):

u r cancelling the stuff of one fraction 45/35 and 20/5 sperately with its own denominator

OpenStudy (jhannybean):

\[\large \frac{8+28v^4}{7v^2}= \frac{8}{7v^2}+\frac{28v^4}{7v^2}\] reduce.

OpenStudy (jhannybean):

Ok. \[\large \frac{4(2+7v^4)}{7v^2}\]

OpenStudy (anonymous):

thats not the answer right ?

OpenStudy (jhannybean):

Your answer choices should contain the variable v^2

OpenStudy (anonymous):

well wats the ans ??

OpenStudy (anonymous):

you guys put adding when i put division

OpenStudy (anonymous):

i put a division sign and when you rewrote it, it had an adding sign

OpenStudy (anonymous):

ooo well sorry .... fine reverse the fraction 20v^3/5v and then cancellation ...

OpenStudy (jhannybean):

Oh, oops. That was a division sign wasnt it...

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

\[\frac{ 40v ^{2} }{ 35v ^{4} }*\frac{ 5v }{20v ^{3} }\]

OpenStudy (jhannybean):

There you go.

OpenStudy (anonymous):

do you cross multiply when you flip them ?

OpenStudy (anonymous):

s

OpenStudy (jhannybean):

You don't have to now, you can cancel.

OpenStudy (anonymous):

so by 5

OpenStudy (anonymous):

s

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

8v^2/7v^4 x v/4v^3

OpenStudy (anonymous):

lol thank you(:

OpenStudy (anonymous):

2/7^4 ... all done

OpenStudy (jhannybean):

good job.